# Quantum amplification and simulation of strong and ultrastrong coupling of light and matter

### v1

@@ 3.3.1 Example: Two atoms simultaneously excited by a single photon
We now briefly discuss a particular example of the simulation of a particular nonlinear process by which two atoms are simultaneously excited by a single photon. This process is a characteristic of ultrastrongly coupled light-matter systems Garziano et al. (2016, 2020); Tomonaga et al. (2023). We consider two atoms coupled to a single cavity mode. The states of the bare basis, working in the dressed-qubit picture, are {|±,±$\{\left|{\pm,\pm,n}\right\rangle\}$. For the nonlinear process where a single photon is simultaneously absorbed by two atoms, the initial and final states are
@@ 3.4 Digital simulation
The simulation protocols described so far in this section all fall into the category of analog quantum simulation, where a system is arranged (in the cases we have seen, by the use of external drives) to behave like the system we want to simulate. The other approach to simulating quantum systems is digital quantum simulation Lloyd (1996); Georgescu et al. (2014), where quantum gates or other means are used to turn different Hamiltonian interactions on and off in a sequence (a Trotter decomposition Trotter (1959)). This yields the same time evolution as the total Hamiltonian we wish to simulate. In examples in this subsection, the proposed and implemented schemes could also be called digital-analog quantum simulations Lamata et al. (2018), where the Hamiltonians available in the simulator are interleaved with quantum gates on parts of the system.
@@ 3.4.1 Theoretical proposal
A proposal for a digital-analog quantum simulation of the quantum Rabi model was first presented in 2014 in Ref. Mezzacapo et al. (2014). Similar to the proposal of Ref. Ballester et al. (2012), discussed in Sec. 3.2.1, the starting point is a system with a qubit coupled to a resonator, which is well described by the JC Hamiltonian in Eq. (127). The key insight used is that the quantum Rabi Hamiltonian
@@
The simulation then consists of short Trotter steps, where the system is switched between evolving in time with the Hamiltonian in Eq. (151) and the Hamiltonian in Eq. (152), by detuning and flipping the qubit. To lowest order in $t/s$, where $t$ is the total evolution time and $s$ is the number of Trotter steps, this approximates the time evolution of the full quantum Rabi Hamiltonian as Trotter (1959)
@@
As is also pointed out in Ref. Mezzacapo et al. (2014), this scheme is readily extended to simulating the Dicke Hamiltonian Dicke (1954)
@@
for $N$ qubits interacting with a single harmonic mode, by using a system well described by the Tavis–Cummings Hamiltonian Tavis and Cummings (1968)
@@
All that is required is to flip and detune all the qubits in parallel in exactly the same way as one qubit is flipped and detuned in the simulation of the quantum Rabi model.
The ideas from Ref. Mezzacapo et al. (2014) were further developed and generalized in Ref. Lamata (2017), which considered simulating some variations of the Dicke model, and also provided more detailed expressions for the error in the Trotter expansion shown in Eq. (153). For example, the Hamiltonian of a Fermi-Bose condensate can be mapped onto a generalized Dicke model with inhomogeneous qubit frequencies Yuzbashyan et al. (2005), which can be simulated just like the standard Dicke Hamiltonian by tuning the qubit frequencies individually. A biased Dicke model, where a term $\sum_{j=1}^{N}\Delta\sigma_{x}^{(j)}$ is added to Eq. (154), appears somewhat more complicated to simulate, requiring more gates applied to the qubits. For a pulsed Dicke model Dasgupta et al. (2015), where the coupling strength $g_{D}$ in Eq. (154) is time-dependent, the simulation scheme of Ref. Mezzacapo et al. (2014) can be modified to include tunable couplings $g$ (provided that the experimental setup allows that).
@@ 3.4.2 Experimental implementation
Figure 13: Digital quantum-simulation protocol implemented in Ref. Langford et al. (2017). The qubit (green line) starts at its sweet spot, detuned from the resonator frequency (red line). At the beginning of a Trotter step, the qubit frequency is tuned close to the resonator frequency. Next, the qubit is detuned again and flipped by a $\pi$ pulse (an additional phase $\phi_{1}$ is introduced by the choice of the rotation axis for the pulse), tuned back close to resonance (to a different frequency than the preceding time), detuned again and flipped back by a second $\pi$ pulse (introducing a second additional phase $\phi_{2}$), tuned close to resonance again (with the same frequency as the first time), and finally detuned back to the sweet spot. The first and last tuning close to resonance implements an effective JC Hamiltonian [Eq. (151)]; the middle tuning close to resonance (with the qubit flipped) implements an effective anti-JC Hamiltonian [Eq. (152)]. The phase difference $\phi_{1}-\phi_{2}$ sets the rotating frame in which the simulation takes place. Figure from Ref. Kockum et al. (2019), A. F. Kockum et al., Nat. Rev. Phys. 1, 19–40 (2019) [adapted from Ref. Langford et al. (2017), N. K. Langford et al., Nat. Commun. 8, 1715 (2017)], with permission.
The proposal of Ref. Mezzacapo et al. (2014) for digital quantum simulation of the quantum Rabi model was implemented experimentally in 2017 Langford et al. (2017), using superconducting circuits. The main components of the setup were a resonator with a resonance frequency $\omega_{\rm cav}/2\pi=6.381\,\mathrm{GHz}$ and a transmon qubit with a tunable frequency (sweet spot for protection from decoherence at $\omega_{q}/2\pi=5.452\,\mathrm{GHz}$). The two were coupled capacitively with a strength $g/2\pi=1.95\,\mathrm{MHz}$, so $g/\omega_{\rm cav}=3\times 10^{-4}$, far from the USC regime. In the simulation, an effective coupling strength up to $g_{R}/\omega_{Rc}\approx 1.8$ was realized.
The experimental protocol is sketched in Fig. 13. To reduce Trotterization error, a second-order Trotter step as described below Eq. (153) was used. Each Trotter step took 122 ns; one simulation run consisted of up to 90 Trotter steps. An improvement on the original protocol in Ref. Mezzacapo et al. (2014) was that the rotation axes used to flip the qubit were chosen to yield phases that defined the frequency of the rotating frame in which the simulation takes place.
@@ 3.5.1 Variational quantum simulation
The broad class of quantum algorithms known as variational quantum algorithms (VQAs) Cerezo et al. (2021) have been widely explored over the past few years, in the hope that they will be able to provide advantages when run on near-term quantum computers still subject to noise. In a VQA, the aim is to solve some problems encoded as minimizing a cost function. To find a solution, an ansatz in the form of a parameterized quantum circuit, applied to some initial quantum states, is chosen. By measuring properties of the output state after the circuit has been applied, information is gathered that lets a classical optimization algorithm update the parameters of the circuit to approach an optimum encoding the solution to the problem. In particular, VQAs can be applied to finding the ground or excited states of a given Hamiltonian in this manner, or even simulating the time evolution; this class of VQAs are known as VQS Yuan et al. (2019); Cerezo et al. (2021).
In 2020, Di Paolo et al. Di Paolo et al. (2020) proposed the use of VQS to find the ground state of a system with an USC. The proposal considered a generalization of the quantum Rabi model [Eq. (126)] to $N$ qubits and $M$ harmonic modes, i.e., a multimode Dicke model, which is described by the Hamiltonian
@@
To treat the problem with VQS, the first step is to encode the bosonic modes in the qubits used in the VQS. The encoding chosen in Ref. Di Paolo et al. (2020) truncated the Fock space of mode $k$ to have at most $n_{k}^{\rm max}$ photons and then encoded this state using $n_{k}^{\rm max}+1$ qubits as $\left|{n_{k}}\right\rangle\rightarrow\left|{0_{0}\ldots 0_{n_{k}-1}1_{n_{k}}0_{n_{k}+1}\ldots 0_{n_{k}^{\rm max}}}\right\rangle$. The annihilation operator of mode $k$ then maps to the qubit operators, such that
@@
and $\boldsymbol{\theta}^{*}$ minimizes the energy of the system [set by the Hamiltonian in Eq. (156)]. In Ref. Di Paolo et al. (2020), the choice of $U(\boldsymbol{\theta})$ is based on the polaron transformation Díaz-Camacho et al. (2016), which approximately disentangles Eq. (156). For one qubit, this transformation is given by
@@
In Ref. Di Paolo et al. (2020), the ability of this variational ansatz to find the ground-state energy of small instances of Eq. (156) was tested both in numerical simulations and in an experiment using three superconducting qubits on an IBM quantum processor. In the numerical simulations, the VQS came within a few percent (relative error) of the ground-state energy for the single- and two-mode quantum Rabi and Dicke models with a Fock-space truncation of 33335555 photons per mode. Interestingly, the energy estimate from the VQS was better than the value given by doing the exact polaron transformation in Eq. (159). In the experiment using superconducting qubits, the errors were larger, but there was qualitative agreement in how the ground-state energy changed as a function of $g/\omega_{q}$ when simulating the resonant quantum Rabi model [Eq. (126)] using two qubits to encode the mode. The errors were attributed to both the noise in the quantum processor and the performance of the classical optimization algorithm used in the VQS.
@@ 3.5.2 Coupled waveguides
A scheme for analog simulation of the dynamics of the quantum Rabi model [Eq. (126)] proposed in 2011 Longhi (2011) takes as its starting point the observation that the state of the system can be written as
@@
The proposal of Ref. Longhi (2011), implemented experimentally in 2012 Crespi et al. (2012), is to construct an array of coupled photonic waveguides where the amplitudes of the propagating light in the waveguides evolve according to Eq. (162). In that setup, the propagation distance in the waveguide corresponds to the time, the coupling $g$ is set by the coupling between adjacent waveguides (which has to increase as $\sqrt{n+1}$ to make $g$ constant; this is achieved by setting the spacing between the waveguides), the qubit transition frequency is set by a modulation of the effective refractive index in the waveguides, and the resonator frequency is determined by the gradient of the refractive index.
In the experiment of Ref. Crespi et al. (2012), this proposal was implemented in 15 waveguides written with a femtosecond laser on a fused silica substrate. The relevant parameters were set such that $g/\omega_{\rm cav}=0.65$, well in the USC regime, was realized. By measuring the light distribution in the waveguides, it was possible to observe the time evolution of the qubit and resonator populations in the simulation. To simulate a different set of parameter values, it was necessary to manufacture a new set of waveguides; no parameters were tunable in situ for this simulator.
@@ 3.5.3 Ultracold atoms in optical lattices
Another proposed method for simulating the physics of the quantum Rabi model uses ultracold atoms in an optical lattice Felicetti et al. (2017a). In the limit of few atoms, the interactions between atoms are negligible and the Hamiltonian of a single atom in the lattice can be written as
@@
where $m$ is the mass of the atom, $p=-i\hbar\frac{\partial}{\partial x}$ is the momentum of the atom, $x$ is the position of the atom, $\omega_{0}$ is the frequency of the atomic motion in a harmonic trap induced by laser driving in the setup, $V$ is the depth of a periodic potential induced by laser driving, and $4k_{0}$ is the wave vector of that potential (stemming from a four-photon interaction with a driving field having a wave vector $k_{0}$ Ritt et al. (2006)).
@@
Rotating the qubit basis appropriately, the Hamiltonian in Eq. (164) corresponds to the quantum Rabi Hamiltonian in Eq. (126) with a cavity frequency $\omega_{\rm cav}=\omega_{0}$, a qubit frequency $\omega_{q}=\frac{V}{2\hbar}$, and a coupling strength $g=2k_{0}\sqrt{\frac{\hbar\omega_{0}}{2m}}$. In Ref. Felicetti et al. (2017a), it is shown that with typical parameters of ultracold rubidium atoms in an optical lattice, these values would correspond to $g/\omega_{\rm cav}\approx 10$, well into the regime of DSC, without much tunability possible, while the effective qubit frequency could be widely tuned in the range $\omega_{q}/\omega_{\rm cav}\in[0,30]$.
The experimental implementation of the proposal from Ref. Felicetti et al. (2017a) was published in 2023 Koch et al. (2023). The experiment employed rubidium atoms in a setup with a trap frequency $\omega_{0}/2\pi$ in the range $[350,750]\>\text{Hz}$ and an effective coupling strength $g/2\pi\in[2290,3090]\>\text{Hz}$, achieving $g/\omega_{\rm cav}$ up to 6.6. The effective qubit frequency $\omega_{q}/2\pi$ could be tuned widely, from 0 Hz to 5850 Hz. These settings enabled the observation of the time evolution of both the bosonic and qubit excitations in simulations of the DSC regime.

### v2

@@ 3.3.1 Example: Two atoms simultaneously excited by a single photon
We now briefly discuss a particular example of the simulation of a particular nonlinear process by which two atoms are simultaneously excited by a single photon. This process is a characteristic of ultrastrongly coupled light-matter systems [171, 175, 174]. We consider two atoms coupled to a single cavity mode. The states of the bare basis, working in the dressed-qubit picture, are {|±,±$\{\left|{\pm,\pm,n}\right\rangle\}$. For the nonlinear process where a single photon is simultaneously absorbed by two atoms, the initial and final states are
@@ 3.4 Digital simulation
The simulation protocols described so far in this section all fall into the category of analog quantum simulation, where a system is arranged (in the cases we have seen, by the use of external drives) to behave like the system we want to simulate. The other approach to simulating quantum systems is digital quantum simulation [176, 178], where quantum gates or other means are used to turn different Hamiltonian interactions on and off in a sequence (a Trotter decomposition [403]). This yields the same time evolution as the total Hamiltonian we wish to simulate. In examples in this subsection, the proposed and implemented schemes could also be called digital-analog quantum simulations [404], where the Hamiltonians available in the simulator are interleaved with quantum gates on parts of the system.
@@ 3.4.1 Theoretical proposal
A proposal for a digital-analog quantum simulation of the quantum Rabi model was first presented in 2014 in Ref. [405]. Similar to the proposal of Ref. [182], discussed in Sec. 3.2.1, the starting point is a system with a qubit coupled to a resonator, which is well described by the JC Hamiltonian in Eq. (130). The key insight used is that the quantum Rabi Hamiltonian
@@
The simulation then consists of short Trotter steps, where the system is switched between evolving in time with the Hamiltonian in Eq. (154) and the Hamiltonian in Eq. (155), by detuning and flipping the qubit. To lowest order in $t/s$, where $t$ is the total evolution time and $s$ is the number of Trotter steps, this approximates the time evolution of the full quantum Rabi Hamiltonian as [403]
@@
As is also pointed out in Ref. [405], this scheme is readily extended to simulating the Dicke Hamiltonian [30]
@@
for $N$ qubits interacting with a single harmonic mode, by using a system well described by the Tavis–Cummings Hamiltonian [31]
@@
All that is required is to flip and detune all the qubits in parallel in exactly the same way as one qubit is flipped and detuned in the simulation of the quantum Rabi model. Note that in order to satisfy the gauge principle, the Dicke Hamiltonian in the dipole gauge is (see Ref. [406] for the case in the Coulomb gauge)
$H_{\rm D}^{\rm dg}=\omega_{\rm cav}a^{\dagger}a+\omega_{q}S_{z}+2ig_{\rm dg}(a^{\dagger}-a)S_{x}+4\eta g_{\rm dg}S_{x}^{2},$ (159)
where $S_{x}=\frac{1}{2}\sum_{j=1}^{N}\sigma_{x}^{(j)}$. It is seen that the Dicke Hamiltonian in Eq. (157) can be viewed as a simpler variant of Eq. (159) in the dipole gauge, with neglecting the diamagnetic term $\propto S_{x}^{2}$. Note that, however, although this diamagnetic term is safely negligible below the USC regime, it can become important in the USC regime [74, 406, 393].
The ideas from Ref. [405] were further developed and generalized in Ref. [407], which considered simulating some variations of the Dicke model, and also provided more detailed expressions for the error in the Trotter expansion shown in Eq. (156). For example, the Hamiltonian of a Fermi-Bose condensate can be mapped onto a generalized Dicke model with inhomogeneous qubit frequencies [408], which can be simulated just like the standard Dicke Hamiltonian by tuning the qubit frequencies individually. A biased Dicke model, where a term $\sum_{j=1}^{N}\Delta\sigma_{x}^{(j)}$ is added to Eq. (157), appears somewhat more complicated to simulate, requiring more gates applied to the qubits. For a pulsed Dicke model [409], where the coupling strength $g_{D}$ in Eq. (157) is time-dependent, the simulation scheme of Ref. [405] can be modified to include tunable couplings $g$ (provided that the experimental setup allows that).
@@ 3.4.2 Experimental implementation
Figure 13: Digital quantum-simulation protocol implemented in Ref. [410]. The qubit (green line) starts at its sweet spot, detuned from the resonator frequency (red line). At the beginning of a Trotter step, the qubit frequency is tuned close to the resonator frequency. Next, the qubit is detuned again and flipped by a $\pi$ pulse (an additional phase $\phi_{1}$ is introduced by the choice of the rotation axis for the pulse), tuned back close to resonance (to a different frequency than the preceding time), detuned again and flipped back by a second $\pi$ pulse (introducing a second additional phase $\phi_{2}$), tuned close to resonance again (with the same frequency as the first time), and finally detuned back to the sweet spot. The first and last tuning close to resonance implements an effective JC Hamiltonian [Eq. (154)]; the middle tuning close to resonance (with the qubit flipped) implements an effective anti-JC Hamiltonian [Eq. (155)]. The phase difference $\phi_{1}-\phi_{2}$ sets the rotating frame in which the simulation takes place. Figure from Ref. [49], A. F. Kockum et al., Nat. Rev. Phys. 1, 19–40 (2019) [adapted from Ref. [410], N. K. Langford et al., Nat. Commun. 8, 1715 (2017)], with permission.
The proposal of Ref. [405] for digital quantum simulation of the quantum Rabi model was implemented experimentally in 2017 [410], using superconducting circuits. The main components of the setup were a resonator with a resonance frequency $\omega_{\rm cav}/2\pi=6.381\,\mathrm{GHz}$ and a transmon qubit with a tunable frequency (sweet spot for protection from decoherence at $\omega_{q}/2\pi=5.452\,\mathrm{GHz}$). The two were coupled capacitively with a strength $g/2\pi=1.95\,\mathrm{MHz}$, so $g/\omega_{\rm cav}=3\times 10^{-4}$, far from the USC regime. In the simulation, an effective coupling strength up to $g_{R}/\omega_{Rc}\approx 1.8$ was realized.
The experimental protocol is sketched in Fig. 13. To reduce Trotterization error, a second-order Trotter step as described below Eq. (156) was used. Each Trotter step took 122 ns; one simulation run consisted of up to 90 Trotter steps. An improvement on the original protocol in Ref. [405] was that the rotation axes used to flip the qubit were chosen to yield phases that defined the frequency of the rotating frame in which the simulation takes place.
@@ 3.5.1 Variational quantum simulation
The broad class of quantum algorithms known as variational quantum algorithms (VQAs) [411] have been widely explored over the past few years, in the hope that they will be able to provide advantages when run on near-term quantum computers still subject to noise. In a VQA, the aim is to solve some problems encoded as minimizing a cost function. To find a solution, an ansatz in the form of a parameterized quantum circuit, applied to some initial quantum states, is chosen. By measuring properties of the output state after the circuit has been applied, information is gathered that lets a classical optimization algorithm update the parameters of the circuit to approach an optimum encoding the solution to the problem. In particular, VQAs can be applied to finding the ground or excited states of a given Hamiltonian in this manner, or even simulating the time evolution; this class of VQAs are known as VQS [412, 411].
In 2020, Di Paolo et al. [413] proposed the use of VQS to find the ground state of a system with an USC. The proposal considered a generalization of the quantum Rabi model [Eq. (126)] to $N$ qubits and $M$ harmonic modes, i.e., a multimode Dicke model, which is described by the Hamiltonian
@@
To treat the problem with VQS, the first step is to encode the bosonic modes in the qubits used in the VQS. The encoding chosen in Ref. [413] truncated the Fock space of mode $k$ to have at most $n_{k}^{\rm max}$ photons and then encoded this state using $n_{k}^{\rm max}+1$ qubits as $\left|{n_{k}}\right\rangle\rightarrow\left|{0_{0}\ldots 0_{n_{k}-1}1_{n_{k}}0_{n_{k}+1}\ldots 0_{n_{k}^{\rm max}}}\right\rangle$. The annihilation operator of mode $k$ then maps to the qubit operators, such that
@@
and $\boldsymbol{\theta}^{*}$ minimizes the energy of the system [set by the Hamiltonian in Eq. (160)]. In Ref. [413], the choice of $U(\boldsymbol{\theta})$ is based on the polaron transformation [414], which approximately disentangles Eq. (160). For one qubit, this transformation is given by
@@
In Ref. [413], the ability of this variational ansatz to find the ground-state energy of small instances of Eq. (160) was tested both in numerical simulations and in an experiment using three superconducting qubits on an IBM quantum processor. In the numerical simulations, the VQS came within a few percent (relative error) of the ground-state energy for the single- and two-mode quantum Rabi and Dicke models with a Fock-space truncation of 33335555 photons per mode. Interestingly, the energy estimate from the VQS was better than the value given by doing the exact polaron transformation in Eq. (163). In the experiment using superconducting qubits, the errors were larger, but there was qualitative agreement in how the ground-state energy changed as a function of $g/\omega_{q}$ when simulating the resonant quantum Rabi model [Eq. (126)] using two qubits to encode the mode. The errors were attributed to both the noise in the quantum processor and the performance of the classical optimization algorithm used in the VQS.
@@ 3.5.2 Coupled waveguides
A scheme for analog simulation of the dynamics of the quantum Rabi model [Eq. (126)] proposed in 2011 [415] takes as its starting point the observation that the state of the system can be written as
@@
The proposal of Ref. [415], implemented experimentally in 2012 [416], is to construct an array of coupled photonic waveguides where the amplitudes of the propagating light in the waveguides evolve according to Eq. (166). In that setup, the propagation distance in the waveguide corresponds to the time, the coupling $g$ is set by the coupling between adjacent waveguides (which has to increase as $\sqrt{n+1}$ to make $g$ constant; this is achieved by setting the spacing between the waveguides), the qubit transition frequency is set by a modulation of the effective refractive index in the waveguides, and the resonator frequency is determined by the gradient of the refractive index.
In the experiment of Ref. [416], this proposal was implemented in 15 waveguides written with a femtosecond laser on a fused silica substrate. The relevant parameters were set such that $g/\omega_{\rm cav}=0.65$, well in the USC regime, was realized. By measuring the light distribution in the waveguides, it was possible to observe the time evolution of the qubit and resonator populations in the simulation. To simulate a different set of parameter values, it was necessary to manufacture a new set of waveguides; no parameters were tunable in situ for this simulator.
@@ 3.5.3 Ultracold atoms in optical lattices
Another proposed method for simulating the physics of the quantum Rabi model uses ultracold atoms in an optical lattice [417]. In the limit of few atoms, the interactions between atoms are negligible and the Hamiltonian of a single atom in the lattice can be written as
@@
where $m$ is the mass of the atom, $p=-i\hbar\frac{\partial}{\partial x}$ is the momentum of the atom, $x$ is the position of the atom, $\omega_{0}$ is the frequency of the atomic motion in a harmonic trap induced by laser driving in the setup, $V$ is the depth of a periodic potential induced by laser driving, and $4k_{0}$ is the wave vector of that potential (stemming from a four-photon interaction with a driving field having a wave vector $k_{0}$ [418]).
@@
Rotating the qubit basis appropriately, the Hamiltonian in Eq. (168) corresponds to the quantum Rabi Hamiltonian in Eq. (126) with a cavity frequency $\omega_{\rm cav}=\omega_{0}$, a qubit frequency $\omega_{q}=\frac{V}{2\hbar}$, and a coupling strength $g=2k_{0}\sqrt{\frac{\hbar\omega_{0}}{2m}}$. In Ref. [417], it is shown that with typical parameters of ultracold rubidium atoms in an optical lattice, these values would correspond to $g/\omega_{\rm cav}\approx 10$, well into the regime of DSC, without much tunability possible, while the effective qubit frequency could be widely tuned in the range $\omega_{q}/\omega_{\rm cav}\in[0,30]$.
The experimental implementation of the proposal from Ref. [417] was published in 2023 [419]. The experiment employed rubidium atoms in a setup with a trap frequency $\omega_{0}/2\pi$ in the range $[350,750]\>\text{Hz}$ and an effective coupling strength $g/2\pi\in[2290,3090]\>\text{Hz}$, achieving $g/\omega_{\rm cav}$ up to 6.6. The effective qubit frequency $\omega_{q}/2\pi$ could be tuned widely, from 0 Hz to 5850 Hz. These settings enabled the observation of the time evolution of both the bosonic and qubit excitations in simulations of the DSC regime. In a follow-up experiment with the same setup, collapse and revival of the excitation number in the DSC regime was observed [420].